A series circuit contains a resistor with R = 40
, an inductor with L = 2 H, a capacitor with
0:0025 F, and a 12
1
Expert's answer
2020-10-25T19:45:33-0400

If 12 is the voltage and the frequency is standard 50 Hz, the magnitude of a current will be


"I=\frac{V}{Z}."

Find the impedance:


"Z=\sqrt{R^2+(X_L-X_C)^2}=\\\space\\=\sqrt{R^2+\bigg(\omega L-\frac{1}{\omega C}\bigg)^2},\\\space\\
I=\frac{V}{\sqrt{R^2+\bigg(\omega L-\frac{1}{\omega C}\bigg)^2}}=0.019\text{ A}."